OpenStudy (anonymous):

Find the standard equation of a circle for Radius 8, Center (3, -5)

OpenStudy (amistre64):

which is standard?

OpenStudy (amistre64):

im thinking its the: (x-Cx)^2 + (y-Cy)^2 = r^2

OpenStudy (amistre64):

where Center = (Cx,Cy)

OpenStudy (anonymous):

I don't know- says find the standard equation of a circle that satisfies the conditions.

OpenStudy (anonymous):

(x -3)^2 + (y+5)^2 = 8^2 x^2 -6x +9 +y^2 +10y +25 = 64 x^2 +y^2 -6x +10y -30 = 0

OpenStudy (amistre64):

in other words, it wants you to know what the standard equation from looks like

OpenStudy (amistre64):

i cant recall if its the tight one or the loose one

OpenStudy (anonymous):

OK- Thanks! Looks like Ritas response might be the one.

OpenStudy (amistre64):

could be :) thats the loose form, its got a few extra pieces hanging off it

OpenStudy (anonymous):

is the tight one easier to follow?